the children to adult ratio is 7 to 2 and if there is 108 people in the zoo how many children are there
Answer by stanbon(75887)

the children to adult ratio is 7 to 2 and if there is 108 people in the zoo how many children are there
----
Equations:
7x + 2x = 108
9x = 108
x = 12
-----
7x = # of children = 7*12 = 84
------
2x = # of adults = 2x = 24
===================
Cheers,
Stan H.
===================